# KEN KANGDA WREN1

### v4

@@ Abstract
Existing approaches to asset-pricing under model-uncertainty adapt classical utility-maximisation frameworks and seek theoretical comprehensiveness. We move toward practice by considering binary model-uncertainties and by switching attention from ’preference’ to ’constraints’. This decomposes economic asset-pricing into the viable pricing of model-risk and non-model risk separately such that the former has a unique and intuitive formula with convenient properties. Its parameter, dynamically conserved under model-risk inference, allows an integrated representation of ex-ante risk-pricing and bias, with ex-post price-effects that can be disentangled, through well-known anomalies, Momentum and Low-Risk, whose risk-reward curves acquire a new significance: peak-rewards measure ex-ante risk-pricing, and peak-locations, bias.
@@ Introduction
Existing works adapting classical theories to incorporate model-risk come under ambiguity-aversion or robust-control, where agents perform recursive optimisation, against goals such as max-min utility, optimal smooth-ambiguity or variational preference, over the set of model choices (e.g. Ju and Miao (2012) or Machina and Siniscalchi (2014)). Related is parameter learning (e.g. Guidolin and Timmermann (2007), Collin-Dufresne et al. (2016)), carried out usually under a definitive model-risk law, without uncertainty.
Existing focus is the nature of preference/utility under model-risk and model-risk discount in a general equilibrium. Our focus is how model-risk pricing may be achieved or characterised ’minimally’, with fewest assumptions and parameters. A natural start is thus the Risk-Neutral Equivalent (RNE) formulation, given the First Theorem of Asset Pricing (FTAP) [^1] Early max-min utility struggled with FTAP-compliance (e.g. Epstein and Wang (1995)). Its modern form (e.g. Chen and Epstein (2002)) and other approaches have no such issues (e.g. Ju and Miao (2012)), especially when reformulated under the banner of smooth-ambiguity (Proposition 3.6 of Burzoni et al. (2021))..
However, viability, No Arbitrage in the FTAP sense, is too weak for most pricing tasks: there are far more viable processes than economically valid ones. Moreover, learning about model-risk makes RNE probabilities path-dependent, and so intractable, even in simple cases (proved formally here, but already observed in Guidolin and Timmermann (2007)).
One way to address such obstacles is to add incremental conditions to the FTAP baseline till RNE formulations that do provide economic asset-pricing can be characterised. It inverts the routine of ’theory making followed by viability checking’, and switches attention from ’agent preference’ to ’structural constraints’. It culminates in the decomposition of viable economic asset-pricing into separately viable components and the identification of a unique and intuitive RNE model-risk pricing formula, whose sole parameter, a constant of motion of the model-risk inferential dynamics, allows an integrated treatment of ex-ante risk-premia and bias such that their ex-post price-effects, only jointly observable, can be disentangled.
@@
We take the existence of representative agents/beliefs for granted, and use the language of rational (Bayesian) inference to mean no more than adherence to the laws of probabilities; likewise, ’rational’ and ’expectation’ signify the same without implying ’objective correctness’, unlike in classical Rational Expectation (RE) frameworks. Three types of probabilities are thus involved: 1) objective ones governing events; 2) subjective ones representing inferential beliefs about the objective world; 3) RNE ones for pricing.
Let our model-risk $B$, of micro or macro economic origin, be binary, with outcome $b$ or $\overline{b}\not=b$, and our $B$-sure models, elementary in the sense of being model-risk free. Such $B$-sure processes may represent the familiar, which by classical RE principles must be well-modelled already. In our context therefore, only change, breaks from routines, may create model-risk (e.g. new CEO, R&D results, or regulatory/governmental/geopolitical events). Many real-world asset valuation challenges fall into this category. Generalisation to model-risks that are non-binary is outlined towards the end, by involving composite $B$-sure processes, which contain discounts for other model-risks.
@@
Our model-risk shall be considered ’one-off’. How it ’flares up’ and the pricing of ’flare-risk’ remains implicit; that is, we study episodic ’model crises’, one at a time. This contrasts with the workhorse Markov regime-switching setting of the literature, where the pricing of the entire process is treated. Further, despite possible uncertainty and indefinite priors, our model-risk is resolvable via regular inference, albeit not before any finite horizon; this is also the case in Hansen-Sargent’s robust-control setting (e.g. Hansen et al. (2006)). Correspondingly, let horizons be sufficiently long for the meaningful development of model-risk inference, and let asset-pricing be time-homogeneous (’identical risk-levels identically priced’) during such developments (i.e. while the model-risk remains active and relevant).
The law of our model-risk, the unconditional distribution of $B$-outcomes, may be unknown. If the true unconditional probability, denoted $\mathbf{p}^{B}_{0}$ throughout, is known, model-detection (learning) and asset-pricing remain classical, with familiar conclusions. If unknown, the same may involve non-classical cognitive and economic elements. Nonetheless, asset-pricing may be formulated in terms of classical inference, with non-classical aspects reflected through preference and discount mechanisms; this is the usual, and our, strategy.
Lastly, data about model-risks (e.g. weekly/monthly indicators of economic regimes) tend to be more frequent and impactful than data about ’routine’ risks (e.g. biannual dividend). Such structures, absent in existing studies but relevant to practice, are treated explicitly here.
@@
Existing Theories in Our Context
@@
Call the set $\mathcal{L}_{\mathcal{S}}:=\{\pi^{I}_{0}Q_{I}|\pi^{I}_{0}\in\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S})),\text{ }I\in\mathcal{I}_{1}(\mathcal{S})\}$ of reference beliefs about the law governing total state space $\mathcal{I}_{1}(\mathcal{S})\times\mathcal{S}$ the set of model-uncertainty, for later reference.
The max-min utility approach offers the simplest strategy: choose at any time the model with the worst outlook from the set $\mathcal{I}_{1}(\mathcal{S})$ of alternatives (in our setting it means pricing for the pessimistic $B$-outcome till $B$-risk resolves). It is a limit of the graded optimisations adopted by the other approaches, where pessimism is tempered by reference beliefs (e.g. Hansen and Sargent (2011) penalises distorted beliefs by their relative entropy to the reference).
Key to all is then the reference belief, by which model-risk pricing is defined and computed. In RE and some parameter-learning settings reference beliefs are known to be true, leading to classical risk-discounts. Under model-uncertainty, even if the reference beliefs happen to be true objectively, perceived uncertainty still triggers extra discounts. All worked examples in the uncertainty literature presume the reference beliefs used by their agents to be unknowingly objectively true; this in our context is unjustifiable.
@@
Any assertion of equality would mask a basic issue facing all attempts to free classical RE frameworks from any built-in coincidence of beliefs and true laws: pricing theory outputs are ex-ante but what can be observed is ex-post, so tests and calibrations are impossible without additional theories of bias. Manifestations of this difficulty include ’observational equivalence’ (’is the large ex-post risk-premium due to bias after all?’; see e.g. Cecchetti et al. (2000)) and ’joint-hypothesis’ (’is an anomaly a sign of bias in the market or in one’s pricing theory?’; see e.g. Fama (2014)). We will illustrate how this issue resolves itself in our approach.
Finally, existing theories require as inputs the total risk as a whole, model and non-model risks together, aggregated across all the assets in a given market. This makes them unwieldy for bottom-up or cross-sectional enquiries, especially when the model-risk(s) can be micro, macro or of a variety of causes. Our method, with its isolation of model-risk pricing and RNE formulation, may prove more flexible and convenient in this context. It also reveals that unless a model-risk is ’consistent’, no interpretable economic model-risk pricing is possible.
@@
Section 1 sets out the usual asset and data structure in continuous time, highlighting the effect of model-risk on standard risk-pricing considerations; Section 2 derives RNE asset-pricing under model-risk; Section 3 demonstrates a potential implication, followed by conclusions and discussions; technical backgrounds and proofs are collected in the Appendix.
It transpires that the RNE formulations of asset-pricing under model-risk are constrained by informational properties (Definition 1, Lemma 1 and Remark 14) and so must follow a canonical form (Proposition 1), with familiar properties (Corollary 1) and price dynamics (Section 2.2.2). One consequence is a potential link between Status Quo Bias and well-known market-anomalies, whose characteristics separate prospective (ex-ante) market risk-pricing from bias (if any) under model-uncertainty, revealing them separately (Section 3.5.1&3.5.2).
@@ 1 Setup
1.1 Usual-Case Assets and Information Structure
For clarity and to avoid clutter, consider assets with a single payoff, some bullet-payoff $Y\in\mathbb{R}$ at horizon (or maturity) $T$, with $1\ll T\leq\infty$; let value and pricing be in logarithm, and time- or risk-free discount be exogenous and nil (with a suitable choice of numéraire). Further:
@@ 1.1 Usual-Case Assets and Information Structure
    Any incremental change to bullet-payoff $Y$, if at all, is public, such that the cumulative firmed-up (’cum-div’) component $\{Z_{t}\}$ of asset-worth, with $Z_{T}\equiv Y$ by definition, has the equation of motion below, subject to binary outcome $B\in\mathcal{B}:=\{+,-\}$, but otherwise based on standard Wiener process $\{w^{Z}_{t}\}$, of natural, usual, filtration $\{\mathcal{F}^{Z}_{t}\}$,
@@
    with $\{\mathcal{F}^{Z}_{t}\}$-adapted volatility $\{\sigma^{Z}_{t}\}$ and drift $\{r^{Z}_{B,t}\}$, on probability space $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$, which is the restriction to $(0,T]$ of the usual Wiener space/measure $(\mathcal{S}^{Z},W^{Z})$ generated by $\{w^{Z}_{t}\}$, with the $B$-sure law ${W}^{Z}_{B}$ of $\{Z_{t}(B)\}$ defined via $d(W^{Z}_{-})_{t}=r^{Z}_{-,t}+\sigma^{Z}_{t}dw^{Z}_{t}$ for $B=-$, and $d(W^{Z}_{+})_{t}=r^{Z}_{+,t}dt+\sigma^{Z}_{t}dw^{Z}_{t}$ for $B=+$; write $r^{Z\Delta}_{t}:=r^{Z}_{+,t}-r^{Z}_{-,t}$.
@@
    There may be an additional $B$-informative real-vector dataflow $\{{D}_{t}(B)\}$, of similar yet $B$-conditionally independent dynamics to $\{Z_{t}(B)\}$:
    $dD_{t}(B)=r^{D}_{B,t}dt+\sigma^{D}_{t}dw^{D}_{t},$ (2)
    all as in (1) but with ’$D$’ labelling; its independence makes it price-irrelevant without model-risk (i.e. under known $B$). Combined dataflow $\{\mathbf{D}_{t}\}:=\{(Z_{t},D_{t})\}$ has state space $\mathcal{S}^{T}:=(\mathcal{S}^{Z}\times\mathcal{S}^{D})|_{T}$ and natural filtration $\{\mathbf{F}_{t}\}:=\{\mathcal{F}^{Z}_{t}\}\vee\{\mathcal{F}^{D}_{t}\}$, under product $B$-sure law $\mathbf{W}^{T}_{B}:=\mathbf{W}_{B}|_{T}$, where $\mathbf{W}_{B}:=({W}^{Z}_{B}\times{W}^{D}_{B})$.
@@
    Current treatments ignore the make up of data $\{\mathbf{D}_{t}\}$, as if only those relevant to model-sure pricing, $\{Z_{t}\}$, exist/matter. We consider situations with data that are exclusive to model-risk (price-irrelevant once model-sure), including cases with trivial model-sure pricing data $\{Z_{t}\}$. It tackles a frequent challenge to partitioners: a crisis market unsettled by a bout of model-risk whose economics and inferential status dominate, while non-model, routine, risks remain, independently evolving or unchanging.
  3. 3.
    Given the above, continuous (predictable) inferential testing of model-risk $B$ takes place on filtered space $(\mathcal{S}^{T}_{\mathcal{B}}:=\mathcal{B}\times\mathcal{S}^{T},\text{ }\{\mathbf{F}_{t}\},\text{ }\mathbf{F}_{T};\text{ }{\pi}^{B(\cdot)}_{0}\mathbf{W}^{T}_{B(\cdot)}(\cdot))$, where ${\pi}^{B(\omega_{B})}_{0}$ is the unconditional belief in outcome $B(\omega_{B})\in\mathcal{B}$, $\omega_{B}:=(B,\omega)\in\mathcal{S}^{T}_{\mathcal{B}}$. Under the condition of $\mathbf{W}_{+}\perp\mathbf{W}_{-}$ with $\mathbf{W}_{+}|_{t}\sim\mathbf{W}_{-}|_{t}$, $\forall t<\infty$, standard testing thus resolves $B$-outcomes at guaranteed resolution-time $T(B)=\inf\big{\{}t|\mathbf{F}_{t}\ni\{B=\cdot\}\big{\}}=\infty$.
    Remark 4.
    The sequential testing of Wiener processes differing by a fixed drift is well-known. Our setup allows time-variations via absolutely continuous clock-change, making finite resolutions ($T(B)<\infty$) possible. Such issues are largely moot under economic risk-pricing (Remark 1), with $T<T(B)\leq\infty$ (see Appendix A for more).
@@
    Such testing, optimal in the sense of Neyman-Pearson Lemma, is based on likelihood-ratios (LR) $\{L^{+/-}_{t}:=\frac{d\mathbf{W}^{T}_{+}|_{t}}{d\mathbf{W}^{T}_{-}|_{t}}\}$, or equivalently, log-LR $\{l^{+/-}_{t}:=\log L^{+/-}_{t}\}$, which form Wiener processes under broad conditions; model-risk $B$ resolves iff. $\lim_{t\to T(B)}l^{+/-}_{t}=(-1)^{\mathbf{1}_{-}(B)}\cdot\infty$ (iff. $\mathbf{W}_{+}|_{T(B)}\perp\mathbf{W}_{-}|_{T(B)}$) for some $T(B)\leq\infty$. Further, standard testing is performed under the equivalent martingale measure of one of the processes (e.g that of $B=-$); henceforth let $\mathbf{W}_{-}$ be by default this base measure, under which $r^{Z}_{-,t}=0=r^{D}_{-,t}$, $\forall t$.
@@
    The above has Bayesian representations: any $L^{+/-}_{t}\in(0,\infty)$ maps bijectively and smoothly to an a posteriori belief ${\pi}^{B}_{t}\in(0,1)$. Bayes’ Rule, in terms of odds-for, $O_{f}[{\pi}^{B}_{t}]:={\pi}^{B}_{t}/\underline{\pi}^{B}_{t}$, with ()¯$\underline{(\cdot)}=1-(\cdot)$, in effect says ’new-odds $=$ old-odds $\times$ likelihood-ratios of data’:
@@
    given a priori odds $O_{f}[{\pi}^{+}_{0}]\in(0,\infty)$; so inferential odds are geometric Wiener processes.
  6. 6.
    The time-$t$, $B$-sure, expected payoff $Y^{B}_{t}:={\mathbf{E}}_{B}[Y|\mathbf{F}_{t}]$ under its $B$-sure law reads ((1)):
    $Y^{B}_{t}=z_{t}+\mathbf{E}_{B}[Y-z_{t}|{\mathbf{F}}_{t}]=z_{t}+\mathbf{E}_{B}[Y-z_{t}]=z_{t}+y_{B}(T-t,t),$ (4)

### v6

@@ Abstract
Existing approaches to asset-pricing under model-uncertainty adapt classical utility-maximisation frameworks and seek theoretical comprehensiveness. We move toward practice by considering binary model-risks and by emphasizing ’constraints’ over ’preference’. This decomposes viable economic asset-pricing into that of model and non-model risks separately, leading to a unique and convenient model-risk pricing formula. Its parameter, a dynamically conserved constant of model-risk inference, allows an integrated representation of ex-ante risk-pricing and bias such that their ex-post impacts are disentangled via well-known anomalies, Momentum and Low-Risk, whose risk-reward patterns acquire a fresh significance: peak-reward reveals ex-ante risk-premia, and peak-location, bias.
@@ Introduction
Existing works adapting classical theories to incorporate model-risk come under the heading of ambiguity-aversion or robust-control, where agents perform recursive optimisation, against goals such as max-min utility, optimal smooth-ambiguity or variational preference, over the set of model choices (e.g. Ju and Miao (2012) or Machina and Siniscalchi (2014)). Related also is parameter learning (e.g. Guidolin and Timmermann (2007), Collin-Dufresne et al. (2016)), often performed under some known model-risk law.
Existing focus is the nature of preference/utility under model-risk and the resulting general equilibrium. We focus here instead on how model-risk pricing might be done or characterised ’minimally’, in terms of assumptions and parameters. A natural start is thus the Risk-Neutral Equivalent (RNE) formulation, given the First Theorem of Asset Pricing (FTAP) [^1] Early max-min utility struggled with FTAP-compliance (e.g. Epstein and Wang (1995)). Its modern form (e.g. Chen and Epstein (2002)) and other approaches have no such issues (e.g. Ju and Miao (2012)), especially when reformulated under the banner of smooth-ambiguity (Proposition 3.6 of Burzoni et al. (2021))..
The RNE approach to model-risk has been seldom discussed, for two good reasons. Firstly, viability, No Arbitrage in the FTAP sense, is too weak for most purposes: there are far more viable price processes than economically valid ones. Secondly, parameter learning makes RNE probabilities path-dependent and intractable, even in simple cases, as observed by Guidolin and Timmermann (2007). We overcome these difficulties by tracing their origins and examining the consequences of the constraints they impose on asset-pricing.
The key turns out to be whether model-risk pricing is a standard Ito process of the model-risk inference process, or equivalently, whether the economic price-of-model-risk has an identical RNE form. It leads to a unique and intuitive model-risk pricing formula, whose parameter, a constant of motion of model-risk inference, allows an integrated treatment of ex-ante risk-premia and bias such that their ex-post effects, only jointly observable, may be disentangled. The RNE approach can apply at the aggregate or the individual-asset level, thereby potentially expanding the data universe helpful to and addressable by model-risk/parameter-learning studies, which are currently closed to stock-level or cross-sectional data.
@@
The existence of representative agents/beliefs is taken for granted, so the language of rational (Bayesian) inference is used, to mean no more than adherence to the laws of probabilities; likewise, ’rational’ and ’expectation’ signify the same without implying ’objective truthfulness’, unlike in classical Rational Expectation (RE) frameworks. We thus work with three types of probabilities: 1) the objectively true ones governing events; 2) the subjective ones representing inferential beliefs about the objective world; and 3) the RNE ones for pricing.
Let our model-risk $B$ have two outcomes, $b$ or $\overline{b}\not=b$, and our $B$-sure models, elementary (’model-risk free’). Such $B$-sure processes represent ’routines’ that by classical RE principles are well-modelled; only change creates model-risk (e.g. new CEO, R&D results, macroeconomic, regulatory or geopolitical events). Non-binary situations that can be handled within the binary framework emerge quite naturally, as will be demonstrated in an application addressing typical questions encountered in practice.
@@
Our model-risk shall be considered one-off. How it ’flares up’ and the pricing of ’flare-risk’ remains implicit; that is, we study episodic ’model crises’, one at a time. This contrasts with the workhorse Markov regime-switching setting of the literature, where the pricing of the entire process is treated. Further, despite possible uncertainty and indefinite priors, our model-risk is resolvable via regular inference, albeit not before any finite horizon; this is also the case in Hansen-Sargent’s robust-control setting (e.g. Hansen et al. (2006)). Correspondingly, let horizons be sufficiently long for the meaningful development of model-risk inference, and let asset-pricing be time-homogeneous (’identical risk-levels identically priced’) during such developments (i.e. while the given model-risk remains active and relevant).
The law of our model-risk, the unconditional distribution of $B$-outcomes, may be unknown. If the objectively true unconditional probability, denoted $\mathbf{p}^{B}_{0}$ throughout, is known, model-detection (learning) and asset-pricing remain classical, with familiar conclusions. If unknown, the same may involve non-classical cognitive and economic elements. Nonetheless, asset-pricing may be formulated in terms of classical inference, with non-classical aspects reflected through preference and discount mechanisms; this is the usual, and our, strategy.
Lastly, data on model-risk (e.g. weekly/monthly indicators of economic regimes) tend to be more frequent and impactful than data on ’routine’ risks (e.g. biannual dividend). Such structures, absent in existing studies but relevant to practice, are treated explicitly here.
@@
From the Perspectives of Standard Model-Risk Studies
@@
The set $\mathcal{L}_{\mathcal{S}}:=\{\pi^{I}_{0}Q_{I}|\pi^{I}_{0}\in\mathcal{I}_{2}(\mathcal{I}_{1}(\mathcal{S})),\text{ }I\in\mathcal{I}_{1}(\mathcal{S})\}$ of reference beliefs about the law of total state space $\mathcal{I}_{1}(\mathcal{S})\times\mathcal{S}$ may be called ’the set of model-uncertainty’, for convenience.
The max-min utility approach offers the simplest solution: choose at any time the model with the worst outlook from the set $\mathcal{I}_{1}(\mathcal{S})$ of alternatives (meaning in our setting choosing the pessimistic $B$-outcome till $B$-risk resolves). It is a limit of the graded optimisations used by others, where pessimism is tempered by reference beliefs (e.g. Hansen and Sargent (2011) penalises distorted beliefs by their relative entropy to a chosen benchmark).
Key to all is then the reference belief, by which model-risk pricing is defined and computed. In RE and some parameter-learning settings reference beliefs are known to be true, leading to classical risk-discounts. Under model-uncertainty, even if the reference beliefs happen to be true objectively, perceived uncertainty still triggers extra discounts. All worked examples in the ambiguity-aversion/robust-control literature presume the reference beliefs of their agents to be unknowingly objectively true; this in our context is hard to justify.
@@
Any assertion of equality would mask a basic issue facing all attempts to free classical RE frameworks from any built-in coincidence of beliefs and true laws: pricing theory outputs are ex-ante but what can be observed is ex-post, so tests and calibrations are impossible without additional theories of bias. Manifestations of this ’impasse’ include ’observational equivalence’ (’is the high realised risk-premium observed due to bias after all?’; see e.g. Cecchetti et al. (2000)) and ’joint-hypothesis’ (’is an anomaly a sign of bias in the market or in one’s pricing theory?’; see e.g. Fama (2014)). We show how this issue may resolve itself in our approach.
Finally, existing theories require ’total risk’ as a whole as inputs, model and non-model risks together, aggregated across all relevant assets in a given market. They are unwieldy for bottom-up (individual-asset) or cross-sectional enquiries, especially when the model-risk(s) can be micro-economic or heterogeneous in origin. Our method, with its isolation of model-risk pricing and RNE formulation, is more flexible and convenient for such applications.
@@
Section 1 sets out the usual asset and data structure in continuous time, highlighting the effect of model-risk on standard risk-pricing considerations; Section 2 derives RNE asset-pricing under model-risk; Section 3 demonstrates a potential application, followed by conclusions and discussions; technical backgrounds and proofs are collected in the Appendix.
We find that the RNE formulations of asset-pricing under model-risk are constrained by economic, informational, necessities (Definition 1 and Lemma 1) and so have a canonical form (Proposition 1), resulting in familiar properties (Corollary 1) and dynamics (Section 2.2.2). One implication is a potential link between Status Quo Bias and well-known anomalies, whose characteristic risk-reward curves reveal and separate ex-ante risk-premia from bias under model-uncertainty (Section 3.4.1&3.4.2).
@@ 1 Setup
1.1 The Usual-Case Asset and Information Structure
For clarity, consider an usual-case asset: it has just one, bullet, payoff, $Y\in\mathbb{R}$, at some horizon or maturity $T$ much longer than unit-time, so $1\ll T<\infty$; let value and price be in logarithm, and time- and risk-free discount exogenous and nil (in a suitable numéraire). Further:
@@ 1.1 The Usual-Case Asset and Information Structure
    Any incremental change (if at all) to bullet-payoff $Y$ is such that the cumulative firmed-up (’cum-div’) asset-worth $\{Z_{t}\}$, with $Z_{T}\equiv Y$, follows the dynamics below, based on a standard Wiener process $\{w^{Z}_{t}\}$, with natural filtration $\{\mathcal{F}^{Z}_{t}\}$, subject to binary outcomes $\{B=b\}$ (e.g. ’high’ or ’low’) labelled by $b\in\mathcal{B}:=\{+,-\}$,
@@
    where its $B$-sure drifts $\{r^{Z}_{B,t}\}$ and $B$-independent volatilities $\{\sigma^{Z}_{t}\}$ are both $\{\mathcal{F}^{Z}_{t}\}$-predictable. The underlying $B$-sure probability space $(\mathcal{S}^{Z}|_{T},W^{Z}_{B}|_{T})$ is the restriction to $(0,T]$ of the Wiener space/measure $(\mathcal{S}^{Z},W^{Z}_{B})$ associated with process $\{Z_{t}(B)\}$, that is, $dW^{Z}_{B,t}=dZ_{t}(B)$. The above allows the sequential testing of Wiener processes of differing drifts [^2] Text-book sequential tests (e.g. Peskir and Shirayev (2006)) concern processes of fixed drift and volatility. Our setup allows time-variations achieved through an absolutely continuous clock-change, or equivalently, $\{\mathcal{F}^{Z}_{t}\}$-predictable drifts and volatilities. For more background see Appendix A..
@@
    The economic impact of model-risk $B$ thus derives from drift difference $r^{Z\Delta}_{t}:=r^{Z}_{+,t}-r^{Z}_{-,t}$; let it be consistent: $Y^{\Delta}(t):=\int_{t}^{T}r^{Z\Delta}_{s}ds>0$ almost surely, at any $t<T$. This is intuitive, also necessary for viable and economically valid risk-pricing, as will be seen (Remark 8).
  3. 3.
    There may be a purely $B$-informative real-vector dataflow $\{{D}_{t}(B)\}$ that is $B$-conditionally independent but otherwise similar to $\{Z_{t}(B)\}$, based on a standard Wiener process $\{w^{D}_{t}\}$,
    $dD_{t}(B)=r^{D}_{B,t}dt+\sigma^{D}_{t}dw^{D}_{t}.$ (2)
    It by definition is price-irrelevant in the absence of model-risk (when $B$-sure). Combined, total dataflow $\{\mathbf{D}_{t}\}:=\{(Z_{t},D_{t})\}$ has state space $\mathcal{S}^{T}:=(\mathcal{S}^{Z}\times\mathcal{S}^{D})|_{T}$ and natural filtration $\{\mathbf{F}_{t}\}:=\{\mathcal{F}^{Z}_{t}\}\vee\{\mathcal{F}^{D}_{t}\}$, under the product $B$-sure law $\mathbf{W}^{T}_{B}:=\mathbf{W}_{B}|_{T}$, with $\mathbf{W}_{B}:={W}^{Z}_{B}\times{W}^{D}_{B}$.
@@
    Existing studies ignore potential structures within dataflow ($\{\mathbf{D}_{t}\}$), as if only data relevant to tangible, firmed-up value, $\{Z_{t}\}$, exist/matter. By considering data that are exclusive to model-risk (price-irrelevant once model-sure), including when tangible-value data, $\{Z_{t}\}$, are absent or trivial, we tackle a frequent challenge facing partitioners: a crisis market unsettled by a bout of model-risk whose economics dominate, while non-model, routine, risks remain, independently evolving or unchanging.
@@
[…]
@@
    Tests based on likelihood-ratios (LR) $L^{+/-}_{t}:=\frac{d\mathbf{W}^{T}_{+}|_{t}}{d\mathbf{W}^{T}_{-}|_{t}}$, or log-LR $l^{+/-}_{t}:=\log L^{+/-}_{t}$, $\forall t<T$, are optimal in the sense of Neyman-Pearson Lemma. Model-risk $B$ resolves iff. $\exists T_{\mathcal{B}}\leq\infty$ such that $\lim_{t\to T_{\mathcal{B}}}l^{+/-}_{t}=(-1)^{\mathbf{1}_{-}(B)}\cdot\infty$ (iff. $\mathbf{W}_{+}|_{T_{\mathcal{B}}}\perp\mathbf{W}_{-}|_{T_{\mathcal{B}}}$). Further, such inferential tests are performed under the equivalent martingale measure of one of the processes (e.g that of $B=$-). Henceforth let $\mathbf{W}_{-}$ be this base measure, under which $r^{Z}_{-,t}=0=r^{D}_{-,t}$, $\forall t$.
  6. 6.
    The above has a Bayesian representation: any $L^{+/-}_{t}\in(0,\infty)$ maps bijectively and smoothly to an a posteriori belief ${\pi}^{B}_{t}\in(0,1)$. Bayes’ Rule, in terms of odds-for, $O_{f}[{\pi}^{B}_{t}]:={\pi}^{B}_{t}/\underline{\pi}^{B}_{t}$, with $\underline{(\cdot)}=1-(\cdot)$, says in effect ’new-odds == old-odds $\times$ likelihood-ratios of data’:
@@
    given a priori odds $O_{f}[{\pi}^{+}_{0}]\in(0,\infty)$. That is, inferential odds are exponentiated log-LR processes; the latter, $\{l^{+/-}_{t}\}$, are Wiener processes under broad conditions ((A.2-A.3)).
  7. 7.
    Given $Z_{t}=z_{t}$, $t<T$, the $B$-sure expected payoff $Y^{B}_{t}:={\mathbf{E}}_{B}[Y|\mathbf{F}_{t}]\equiv{\mathbf{E}}_{B}[Y|\mathcal{F}^{Z}_{t}]$ reads ((1-2)):
    $Y^{B}_{t}=z_{t}+\mathbf{E}_{B}[Y-z_{t}|{\mathcal{F}^{Z}_{t}}]=z_{t}+\mathbf{E}_{B}[Y-z_{t}]=z_{t}+y_{B}(T-t,t),$ (4)
    where $y_{B}(T-t,t)-y_{B}(T-t-t^{\prime},t+t^{\prime})=\mathbf{E}_{B}[\int_{t}^{t+t^{\prime}}dZ_{s}(B)]$, $\forall t^{\prime}\leq T-t$, and so $y_{B}(T-t,t)=\int_{t}^{T}{r}^{Z}_{B,s}ds$, as by definition $\{Y^{B}_{t}\}$ is a martingale and $y_{B}(0,T)\equiv 0$. Note also that this relates to model-risk impact $Y^{\Delta}(t)$ via $Y^{\Delta}(t)\equiv Y^{+}_{t}-Y^{-}_{t}\equiv{y}_{+}(T-t,t)-{y}_{-}(T-t,t)$.
1.2 The FTAP Baseline under Model-Risk